# Efficient tomography of microwave photonic cluster states

### v2

@@ IV.1 Estimating the required bond dimension
Figure 3: Estimating the required bond dimension of the MPO representation using the measured local correlations. (a) Schematic of the procedure. (b), (c) Five largest singular values $\sigma_{1},\ldots,\sigma_{5}$ of the four-qubit local correlation matrices $B_{s}$ ($s=1,\ldots,N-3$) measured for the $N$-qubit linear cluster states with $N=10$ and 35, respectively. Error bars represent the standard errors of the singular values estimated by propagating the statistical errors of the correlation data. Because only four singular values of any $B_{s}$ are significantly larger than zero, the bond dimension of these cluster states can be estimated as $D=4$.
The state of a photon chain generated by our photon source has a bond dimension of at most $164^{2}=$ 16 because only four energy levels participate in the photon generation even if there are leakage errors into the fourth energy level of the transmon. However, to reduce the computational cost of the MPO reconstruction, it is desirable to know in advance if the measured state has a more efficient MPO representation with a smaller bond dimension. In fact, the state generated by our experimental protocol can be expected to have a bond dimension of $42^{2}=$ 4 if there is no leaked population outside the qubit subspace $\{\lvert g0\rangle,\lvert e0\rangle\}$ after each conditional photon emission.
@@
Since the measured four-qubit correlations $B_{s}$ contain statistical noise, we estimate $\operatorname{Rank}[B_{s}]$ by counting the number of singular values of $B_{s}$ which are significantly larger than their statistical uncertainties. The statistical uncertainties are estimated by propagating the uncertainties of the quadrature measurements to the correlation matrix $B_{s}$ then to the singular values (see Appendix H for details). Figures 3(b) and (c) show the five largest singular values $\sigma_{1},\ldots,\sigma_{5}$ of the four-qubit correlation matrices $B_{s}$ measured for an $N$-qubit linear cluster state with $N=10$ and 35, respectively. Because only four singular values of any $B_{s}$ are significantly larger than zero, the bond dimension required to express these cluster states can be estimated as $D=4$, up to the statistical uncertainties of the quadrature measurements.
@@ IV.2 Least-squares fitting by an MPO
In this work, the measured Pauli-basis correlations $C^{(a,b,c,d,e)}_{s}$ contain statistical uncertainties of various magnitudes because they are calculated using Eqs. (7a)=̄(7c) from the quadrature moments. In particular, Pauli-basis correlations which measure many of the qubits in the $\hat{Z}_{s}$ basis have large uncertainties because they are calculated from high-order quadrature moments. To avoid overfitting to the large statistical errors in such correlation measurements, we perform a weighted least-squares fitting, where each residual is weighted by the reciprocal of the standard error of the fit target (see Appendix J for the fit result for the 10-qubit cluster state).
@@ IV.3 Results of MPO reconstruction
Figure 5: Quantum state fidelities of the generated $N$-qubit cluster states to the ideal cluster states. (a) Fidelity of the cluster state $\mathcal{F}$ (blue circle) and the fidelities calculated using numerical models with only photon loss and dephasing errors, where the error probabilities are individually fit to each qubit for each $N$ (orange square) or are extrapolated from the average values for $N=5$ (green triangle). (b) Relative uncertainty of $\mathcal{F}$ calculated by propagating the uncertainties of the quadrature measurements.
@@
Figure 5(a) shows the quantum state fidelities of the $N$-qubit cluster states to the ideal cluster states. For comparison, we also plot the fidelities calculated using numerical models which apply the photon loss and dephasing errors to each qubit in the ideal cluster state. The error probabilities used in the numerical models are obtained from the measurements of the mean photon numbers and the stabilizer operators (see Appendix F for details). For $N>10$, the fidelities of the cluster states are significantly smaller than that of the numerical model with 9.8% photon loss and 4.6% phase flip errors, which are the average error probabilities obtained for $N=5$. This suggests that the coherence properties of the photon source used in this work degrade when a large number of photonic qubits are generated in succession, possibly due to heating or quasiparticle generation by the strong pulses driving the $\lvert f0\rangle$$\lvert g1\rangle$ transition. By allowing the error probabilities in the numerical model to differ for each qubit and for each $N$, the fidelities of the cluster states can be more closely explained by the photon loss and dephasing error models. However, the fidelities of the 30- and 35-qubit cluster states still deviate significantly from the numerical model, which suggests that there are error processes which cannot be modeled by a combination of photon loss and dephasing.
Figure 5(b) shows the relative uncertainty of the fidelity calculated by propagating the uncertainties of the quadrature measurements. For the fixed measurement repetition of $10^{7}$ per measurement setting, the relative uncertainties increase only linearly with the length of the cluster state. This demonstrates that the tomography method is efficient in terms of the statistical uncertainty because the number of measurement repetitions required to achieve a given error tolerance can be expected to only increase quadratically with the number of qubits.
@@
IV.4 Lower bound of fidelity based on stabilizer measurements
It is possible to use the set of stabilizer measurements, which are local correlation measurements of up to three qubits, to efficiently obtain a lower bound of the fidelity of the state being measured to an ideal linear cluster state [28]. This method is more general in the sense that it does not assume that the state can be efficiently represented by an MPO. The lower bound for the fidelity $\mathcal{F}$ is given by
$\mathcal{F}\geq\prod_{s=1,3,\ldots}\frac{1+\langle\hat{S}_{s}\rangle}{2}+\prod_{s=2,4,\ldots}\frac{1+\langle\hat{S}_{s}\rangle}{2}-1,$ (10)
where
$\hat{S}_{1}$ $\coloneqq\hat{X}_{1}\hat{Z}_{2},$ (11a)
$\hat{S}_{N}$ $\coloneqq\hat{Z}_{N-1}\hat{X}_{N}$ (11c)
are the stabilizer operators of the $N$-qubit linear cluster state. This lower bound was used in Refs. 17 and 18 to efficiently evaluate the fidelities of their photonic cluster states. However, a lower bound of 0.40 is obtained for the five-qubit cluster state generated in this work, which significantly underestimates the actual fidelity of $\mathcal{F}=0.616\pm 0.006$ obtained from the result of the full tomography detailed in Appendix F. For the larger cluster states, the right-hand side of Eq. (10) becomes negative and does not give a meaningful lower bound. Therefore, the lower bound based on the stabilizer measurements is not effective for evaluating the fidelities of these cluster states. It remains to be seen if the efficient lower bound recently proposed in Ref. 53, which uses additional measurement settings, provides a tighter lower bound of the fidelity.
@@ V.2 Lower bound based on stabilizer measurements
Reference 29 proposed a lower bound for the localizable entanglement of a linear cluster state which can be efficiently obtained using the stabilizer measurements. It uses the localizable entanglement defined in terms of the concurrence $\mathcal{C}_{\mathrm{le}}$, which is a measure of two-qubit entanglement taking the maximum of $\mathcal{C}_{\mathrm{le}}=1$ for a Bell state and the minimum of $\mathcal{C}_{\mathrm{le}}=0$ for a separable state [55, 56]. The lower bound for the localizable entanglement between the $s$th and $(s+k)$th qubits is given by
@@
where $\hat{S}_{r}$ are the stabilizer operators of the linear cluster state given in Eqs. (11a)=̄(11c). This lower bound does not assume that the state has an efficient MPO representation.
@@ VI DISCUSSIONS
Another interesting avenue for future work is to generalize the tomography method to entangled states with more complex graph structures. To efficiently parameterize such states, tensor-network states with various structures, including trees, lattices, and their continuum limits, have been actively studied [66]. In particular, the two-dimensional cluster states which have been generated in the optical and microwave domains can be efficiently represented by a projected entangled-pair operator (PEPO) [67]. A PEPO can be seen as a generalization of an MPO to an arbitrary graph or a generalization of a PEPS to a mixed state. Note that the fact that a two-dimensional cluster state can be efficiently parameterized does not imply that measurement-based quantum computation is classically simulatable because it is still exponentially costly to obtain the result of the computation by contracting the two-dimensional tensor network. Developing an efficient reconstruction algorithm for two-dimensional entangled states requires further studies because our present work makes use of the fact that a one-dimensional tensor network can be efficiently contracted.
@@ ACKNOWLEDGMENTS.
This work was supported in part by the University of Tokyo Program of Excellence in Photon Science (XPS), the Japan Society for the Promotion of Science (JSPS) Fellowship (Grant No. JP22J13650), the Japan Science and Technology Agency (JST) Exploratory Research for Advanced Technology (ERATO) project (Grant No. JPMJER1601), the Ministry of Education, Culture, Sports, Science and Technology (MEXT) Quantum Leap Flagship Program (Q-LEAP) (Grant No. JPMXS0118068682), the JSPS Grant-in-Aid for Scientific Research (KAKENHI) (Grant Nos. JP22H04937, 23H03818, and 22K18682), the JST Center of Innovation for Sustainable Quantum AI (Grant No. JPMJPF2221). T. O. acknowledges the support from the Endowed Project for Quantum Software Research and Education, the University of Tokyo.
@@
$\lvert g\rangle$$\lvert e\rangle$ energy relaxation time $T_{1}$ 14.3$\pm$0.9 $\mu$s
$\lvert g\rangle$$\lvert e\rangle$ Ramsey dephasing time $T_{2}^{*}$ 6.9$\pm$0.7 $\mu$s
$\lvert g\rangle$$\lvert e\rangle$ Hahn-echo dephasing time $T_{2}^{\mathrm{echo}}$ 9.9$\pm$1.4 $\mu$s
$\lvert e\rangle$$\lvert f\rangle$ energy relaxation time $T_{1f}$ 15.1$\pm$1.5 $\mu$s
$\lvert e\rangle$$\lvert f\rangle$ Ramsey dephasing time $T_{2ef}^{*}$ 5.9$\pm$0.6 $\mu$s
@@
The experimental setup used in this work is nominally identical to the one used in Ref. 42 except that the JPA is replaced by a similar one with a larger bandwidth and the 8–12 GHz bandpass filter between the photon source and the JPA is replaced by a 9–11 GHz bandpass filter. The sample for the photon source is also nominally identical except that the resonator is shorter and the gap between the resonator and the coupling pin is larger. Table 1 shows the measured sample parameters.
@@
Here, we introduce the formalism of matrix product operators (MPO), which can efficiently represent a certain class of many-body entangled states [31]. The notation used in this work is inspired by Refs. 34 and 68.
@@ B.1 Matrix product state
The MPO formalism is a generalization of the matrix-product-state (MPS) formalism, which we introduce here.
@@
An MPS representation of this state is obtained by decomposing the tensor $\psi_{i_{1},\ldots,i_{N}}$ into a contraction of $N$ tensors: a $2\times D$ tensor $A_{1}$, $D\times 2\times D$ tensors $A_{2},\cdots,A_{N-1}$, and a $D\times 2$ tensor $A_{N}$. The contraction can be expressed as matrix multiplications by slicing each tensor along the 2-dimensional axis:
@@
The 2-dimensional axis of the tensor $A_{s}$ corresponds to the basis states $\lvert 0\rangle$ and $\lvert 1\rangle$ of the $s$th qubit. For notational convenience, let us define the ket-valued matrices
@@ B.4 Quantum state fidelity
B.5 Local error
@@ C.1 MPS representation
Figure 10: Transformation rules used in Figs. 9(a)=̄(c). (a) A controlled-Z gate equals a Hadamard matrix and two COPY tensors. (b) A |+delimited-|$\lvert$+$\rangle$| + ⟩ state and a COPY tensor equals an identity matrix.
@@
To find an MPS representation of the state generated by this circuit, we interpret the quantum circuit as a tensor network. We then apply the transformation rules shown in Figs. 10(a) and (b) to simplify the tensor network as shown in Figs. 9(b) and (c), respectively (see Ref. 38 for other transformation rules). Here, we have used the Hadamard matrix
@@
$\mathsf{A}_{N}$ $=H\begin{bmatrix}\lvert 0\rangle\\ \lvert 1\rangle\end{bmatrix}=\begin{bmatrix}\lvert+\rangle\\ \lvert-\rangle\end{bmatrix}.$ (32c)
@@ D.3 Generating a linear cluster state
Figure 14: Protocol for generating a photonic linear cluster state implemented by the pulse sequence in Fig. 2(e). (a) Quantum-circuit representation, which can also be interpreted as a tensor network. (b) Tensor network after applying the transformation rules shown in Figs. 15(a)=̄(d), which matches the tensor-network representation of a linear cluster state shown in Fig. 9(c).
Figure 15: Transformation rules used in Fig. 14(b). (a) A $\lvert 0\rangle$ state and a $\pi/2$ rotation equals a |+delimited-|$\lvert$+$\rangle$| + ⟩ state. (b) An $X$ gate and a $\pi/2$ rotation equals a Hadamard gate. (c) A $\lvert 0\rangle$ state and an XOR tensor equals an identity matrix. (d) A SWAP gate equals crossed lines.
@@
is the quantum gate implemented by the $\pi/2$ pulse. By applying the transformation rules shown in Figs. 15(a)=̄(d), one can obtain the tensor-network representation of a linear cluster state shown in Fig. 9(c).

### v3

@@ IV.1 Estimating the required bond dimension
Figure 3: Estimating the required bond dimension of the MPO representation using the measured local correlations. (a) Schematic of the procedure. (b), (c) Five largest singular values $\sigma_{1},\ldots,\sigma_{5}$ of the four-qubit local correlation matrices $B_{s}$ ($s=1,\ldots,N-3$) measured for the $N$-qubit linear cluster states with $N=10$ and 35, respectively. Error bars represent the standard errors of the singular values estimated by the first-order propagation of the statistical errors of the correlation data. Because only four singular values of any $B_{s}$ are significantly larger than zero, the bond dimension of these cluster states can be estimated as $D=4$.
The state of a photon chain generated by our photon source has a bond dimension of at most $4^{2}=16$ because only four energy levels participate in the photon generation even if there are leakage errors into the fourth energy level of the transmon. However, to reduce the computational cost of the MPO reconstruction, it is desirable to know in advance if the measured state has a more efficient MPO representation with a smaller bond dimension. In fact, the state generated by our experimental protocol can be expected to have a bond dimension of $2^{2}=4$ if there is no leaked population outside the qubit subspace $\{\lvert g0\rangle,\lvert e0\rangle\}$ after each conditional photon emission.
@@
Since the measured four-qubit correlations $B_{s}$ contain statistical noise, we estimate $\operatorname{Rank}[B_{s}]$ by counting the number of singular values of $B_{s}$ which are significantly larger than their statistical uncertainties. The statistical uncertainties are estimated by the first-order propagation of the uncertainties of the quadrature measurements to the correlation matrix $B_{s}$ then to the singular values (see Appendix H for details). Figures 3(b) and (c) show the five largest singular values $\sigma_{1},\ldots,\sigma_{5}$ of the four-qubit correlation matrices $B_{s}$ measured for an $N$-qubit linear cluster state with $N=10$ and 35, respectively. Because only four singular values of any $B_{s}$ are significantly larger than zero, the bond dimension required to express these cluster states can be estimated as $D=4$, up to the statistical uncertainties of the quadrature measurements. Note that this is only a qualitative assessment and that a statistical test for the rank of a noisy matrix should be developed in the future to rigorously determine the bond dimension of the measured state.
@@ IV.2 Least-squares fitting by an MPO
In this work, the measured Pauli-basis correlations $C^{(a,b,c,d,e)}_{s}$ contain statistical uncertainties of various magnitudes because they are calculated using Eqs. (7a)(7c) from the quadrature moments. In particular, Pauli-basis correlations which measure many of the qubits in the $\hat{Z}_{s}$ basis have large uncertainties because they are calculated from high-order quadrature moments. To avoid overfitting to the large statistical errors in such correlation measurements, we perform a weighted least-squares fitting, where each residual is weighted by the reciprocal of the standard error of the fit target (see Appendix J for the fit result for the 10-qubit cluster state).
@@ IV.3 Results of MPO reconstruction
Figure 5: Quantum state fidelities and purities of the generated $N$-qubit cluster states. (a) Fidelities $\mathcal{F}$ of the generated cluster states to the ideal cluster states (blue circle). Also plotted are the fidelities calculated using numerical models with only photon loss and dephasing errors, where the error probabilities are individually fit to each qubit for each NN (orange square) or are extrapolated from the average values for $N=5$ (green triangle). (b) Same plot for the purity $\gamma$. (c) Relative uncertainty of $\mathcal{F}$ calculated by propagating the uncertainties of the quadrature measurements. (d) Same plot for the purity $\gamma$.
@@
Figure 5(a) shows the quantum state fidelities of the $N$-qubit cluster states to the ideal cluster states. For comparison, we also plot the fidelities calculated using numerical models which apply the photon loss and dephasing errors to each qubit in the ideal cluster state. The error probabilities used in the numerical models are obtained from the measurements of the mean photon numbers and the stabilizer operators (see Appendix F for details). Figure 5(b) is the same plot for the purity $\gamma\coloneqq\operatorname{Tr}[\hat{\rho}^{2}]$, which takes the maximum of $\gamma=1$ for any pure state and the minimum of $\gamma=2^{-N}$ for the completely mixed state of $N$ qubits. For $N>10$, the fidelities of the cluster states are significantly smaller than that of the numerical model with 9.8% photon loss and 4.6% phase flip errors, which are the average error probabilities obtained for $N=5$. By allowing the error probabilities in the numerical model to differ for each qubit and for each $N$, the fidelities and purities of the cluster states can be closely explained by the photon loss and dephasing error models for $N=15$, 20, and 25. This suggests that the generated cluster states for $N=15$, 20, and 25 are affected by significantly larger incoherent errors than for $N=5$ and 10, possibly due to heating or quasiparticle generation by the strong pulses driving the $\lvert f0\rangle$$\lvert g1\rangle$ transition. However, the fidelities of the 30- and 35-qubit cluster states still deviate significantly from the numerical model, which means that there are coherent error processes in addition to the increased incoherent errors. This suggests that the gate calibrations have drifted significantly between the experiments for $N=25$ and 30.
Figures 5(c) and (d) show the relative uncertainties of the fidelity and the purity, respectively, calculated by propagating the uncertainties of the quadrature measurements. For the fixed measurement repetition of $10^{7}$ per measurement setting, the relative uncertainties increase only linearly with the length of the cluster state. This demonstrates that the tomography method is efficient in terms of the statistical uncertainty because the number of measurement repetitions required to achieve a given error tolerance can be expected to only increase quadratically with the number of qubits.
@@ V.2 Lower bound based on stabilizer measurements
Reference 43 proposed a lower bound for the localizable entanglement of a linear cluster state which can be efficiently obtained using the stabilizer measurements. It uses the localizable entanglement defined in terms of the concurrence $\mathcal{C}_{\mathrm{le}}$, which is a measure of two-qubit entanglement taking the maximum of $\mathcal{C}_{\mathrm{le}}=1$ for a Bell state and the minimum of $\mathcal{C}_{\mathrm{le}}=0$ for a separable state [22, 66]. The lower bound for the localizable entanglement between the $s$th and $(s+k)$th qubits is given by
@@
where
$\hat{S}_{1}$ $\coloneqq\hat{X}_{1}\hat{Z}_{2},$ (11a)
$\hat{S}_{N}$ $\coloneqq\hat{Z}_{N-1}\hat{X}_{N}$ (11c)
are the stabilizer operators of the $N$-qubit linear cluster state. This lower bound does not assume that the state has an efficient MPO representation.
@@ VI DISCUSSIONS
Another interesting avenue for future work is to generalize the tomography method to entangled states with higher-dimensional graph structures. For example, an efficient tomography method for a two-dimensional cluster state will be desired if the generation scheme demonstrated in Ref. 44 is scaled up to use an array of $M\gg 2$ photon emitters and an entangling gate between every neighboring pair. For such a system, our tomography scheme currently does not scale polynomially because the bond dimension $D$ is determined by the collective system dimension of the $M$ photon emitters, which increases exponentially with $M$. In such a case, it becomes necessary to efficiently parameterize the photonic state using the spatial locality of the entanglement among the emitters [42] in addition to the time-domain locality utilized in this work. This is possible using the projected entangled-pair operator (PEPO) representation, which can be seen as a generalization of an MPO to an arbitrary graph or a generalization of a PEPS to a mixed state [18]. Note that the fact that a two-dimensional cluster state can be efficiently parameterized does not imply that measurement-based quantum computation is classically simulatable because it is still exponentially costly to obtain the result of the computation by contracting the two-dimensional tensor network. Developing an efficient reconstruction algorithm for two-dimensional entangled states requires further studies because our present work makes use of the fact that a one-dimensional tensor network can be efficiently contracted.
@@ ACKNOWLEDGMENTS.
We thank Yan Li for pointing out an error in Eq. (10) of the preprint version, which rendered our analysis in Sec. IV.D incorrect. Section IV.D was deleted from this version of the manuscript without compromising our main claims. This work was supported in part by the University of Tokyo Program of Excellence in Photon Science (XPS), the Japan Society for the Promotion of Science (JSPS) Fellowship (Grant No. JP22J13650), the Japan Science and Technology Agency (JST) Exploratory Research for Advanced Technology (ERATO) project (Grant No. JPMJER1601), the Ministry of Education, Culture, Sports, Science and Technology (MEXT) Quantum Leap Flagship Program (Q-LEAP) (Grant No. JPMXS0118068682), the JSPS Grant-in-Aid for Scientific Research (KAKENHI) (Grants No. JP22H04937, No. 23H03818, and No. 22K18682), the JST Center of Innovation for Sustainable Quantum AI (Grant No. JPMJPF2221). T. O. acknowledges the support from the Endowed Project for Quantum Software Research and Education, the University of Tokyo.
@@
[…]
@@
The experimental setup used in this work is nominally identical to the one used in Ref. 56 except that the JPA is replaced by a similar one with a larger bandwidth and the 8–12 GHz bandpass filter between the photon source and the JPA is replaced by a 9–11 GHz bandpass filter. The sample for the photon source is also nominally identical except that the resonator is shorter and the gap between the resonator and the coupling pin is larger. Table 3 shows the measured sample parameters. The frequency, external and internal linewidths, and dispersive shift of the resonator are determined using the post-selected resonator spectroscopy described in Ref. 56, which uses quantum nondemolition readouts to ensure that the qubit stays in either $\lvert g\rangle$ or $\lvert e\rangle$ while the reflection spectrum of the resonator is being measured.
@@
Here, we introduce the formalism of MPOs, which can efficiently represent a certain class of many-body entangled states [17]. The notation used in this work is inspired by Refs. 8 and 9.
@@ B.1 Matrix product state
The MPO formalism is a generalization of the MPS formalism, which we introduce here.
@@
An MPS representation of this state is obtained by decomposing the tensor $\psi_{i_{1},\ldots,i_{N}}$ into a contraction of $N$ tensors: a $2\times D$ tensor $A_{1}$, $D\times 2\times D$ tensors $A_{2},\cdots,A_{N-1}$, and a $D\times 2$ tensor $A_{N}$. The contraction can be expressed as matrix multiplications by slicing each tensor along the two-dimensional axis:
@@
The two-dimensional axis of the tensor $A_{s}$ corresponds to the basis states $\lvert 0\rangle$ and $\lvert 1\rangle$ of the $s$th qubit. For notational convenience, let us define the ket-valued matrices
@@ B.4 Quantum state fidelity
B.5 Purity
The purity of a state $\hat{\rho}$ is defined as
$\gamma\coloneqq\operatorname{Tr}[\hat{\rho}^{2}].$ (23)
In the MPO formalism, it can be calculated as
$\gamma=\frac{1}{2^{N}}\sum_{i_{1},\ldots,i_{N}\in\{0,1,2,3\}}A^{(i_{1})}_{1}\cdots A^{(i_{N})}_{N}A^{(i_{1})}_{1}\cdots A^{(i_{N})}_{N}$. (24)
This equation is represented by the same tensor-network diagram as Fig. 7(a) but with $\hat{\mathsf{T}}_{s}$ replaced by $\hat{\mathsf{A}}_{s}$. The partial derivative of the purity by a matrix element of the MPO is twice the value represented by the tensor-network diagram in Fig. 7(b) with $\hat{\mathsf{T}}_{s}$ replaced by $\hat{\mathsf{A}}_{s}$.
B.6 Local error
@@ C.1 MPS representation
Figure 10: Transformation rules used in Figs. 9(a)(c). (a) A controlled-$Z$ gate equals a Hadamard matrix and two COPY tensors. (b) A $\lvert$$\rangle$ state and a COPY tensor equals an identity matrix.
@@
To find an MPS representation of the state generated by this circuit, we interpret the quantum circuit as a tensor network. We then apply the transformation rules shown in Figs. 10(a) and (b) to simplify the tensor network as shown in Figs. 9(b) and (c), respectively (see Ref. 14 for other transformation rules). Here, we have used the Hadamard matrix
@@
[…]
@@ D.3 Generating a linear cluster state
Figure 14: Protocol for generating a photonic linear cluster state implemented by the pulse sequence in Fig. 2(e). (a) Quantum-circuit representation, which can also be interpreted as a tensor network. (b) Tensor network after applying the transformation rules shown in Figs. 15(a)(d), which matches the tensor-network representation of a linear cluster state shown in Fig. 9(c).
Figure 15: Transformation rules used in Fig. 14(b). (a) A $\lvert 0\rangle$ state and a $\pi/2$ rotation equals a $\lvert$$\rangle$ state. (b) An $X$ gate and a $\pi/2$ rotation equals a Hadamard gate. (c) A $\lvert 0\rangle$ state and an XOR tensor equals an identity matrix. (d) A SWAP gate equals crossed lines.
@@
is the quantum gate implemented by the $\pi/2$ pulse. By applying the transformation rules shown in Figs. 15(a)(d), one can obtain the tensor-network representation of a linear cluster state shown in Fig. 9(c).